A $1,000 par bond with a 4% semi-annual coupon has 10 years to maturity trades at a yield of 6%. What would be it’s price?

  • A. $851.23

  • B. $922.58

  • C. $1,148.77

  • D. $541.20
Answer #1

Answer: Option A is correct.
Face value=$1000
Coupon rate=4%
Semiannual coupon rate=4%/2=2%
Semiannual coupon payment=(Semiannual coupon rate)*(Face value)=2%*1000=20

Time period=10 years
As the coupon payments are made semiannually, the number of periods =10*2=20
Yield to maturity=6%
Semiannual yield to maturity=6%/2=3%
We can determine the present value of the bond using excel.

1 Face value 1000 2 Payment 3 Number of periods 20 4 Yield to maturity 3% 5 Present value= ($851.23) 6 Formula used: PV(14,V3

So, the bond price =$851.23